# Discrete Chi-Square Method can model and forecast complex time series

### v4

@@ Abstract
We show how intensive, large and accurate time series can allow us to see through time. Many phenomena have aperiodic and periodic components. An ideal time series analysis method would detect such trend and signal(-s) combinations. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection. We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The DCM can forecast complex time series. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
@@
Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
@@ 2.1 Discrete Chi-square method (DCM)
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [24, TSPA, Equation 1] and the Continuous Period Search [29, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [21] [^2] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [21, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [25, AL6]
@@
The time series is non-stationary because it contains “a trend” [33, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [25, AL9].
@@ 3.10 Significance estimates
For this Model 1 time series, there is no need to discuss the DFT significance estimates [21, their Equation 22] because this method fails to detect the correct period.
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 20, 22, 26]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@ 4 Discussion
[…]
Acknowledgements This work has made use of NASA’s Astrophysics Data System (ADS).
[…]
Author contribution Lauri Jetsu performed this research and wrote the manuscript.
[…]
Consent for publication The author of this manuscript grants permission for the publication in Journal of Statistical Theory and Practice.
[…]
Data availability All data files and all DCM control files are stored to
[…]
@@
There is also a manual of how to use those codes.

### v5

@@ Abstract
We show how intensive, large and accurate time series can allow us to see through time. Many phenomena have aperiodic and periodic components. An ideal time series analysis method would detect such trend and signal(-s) combinations. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection. We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The DCM can forecast complex time series. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series. In our appendix, we show that DCM can model and forecast El Niño.
@@
Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform, El Niño
@@ 2.1 Discrete Chi-square method (DCM)
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [29, TSPA, Equation 1] and the Continuous Period Search [36, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [23] [^2] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [23, Equation 1]. The notations for the DFT model are
@@
The time series is “too short” [30, AL6]
@@
The time series is non-stationary because it contains “a trend” [42, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [30, AL9].
@@ 3.10 Significance estimates
For this Model 1 time series, there is no need to discuss the DFT significance estimates [23, their Equation 22] because this method fails to detect the correct period.
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 22, 26, 31]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@ 4 Discussion
5 Appendix
In this appendix, we show that our DCM can model and forecast the El Niño phenomenon between the years 1870 and 2024. This result may be the “holy grail” of climatology. The solar forcing most probably causes the detected $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ year signals. These strict periodicities indicate that the solar cycle is caused by the planets. Our results confirm that the DCM is an ideal time series analysis method for detecting an unknown number of signals superimposed on an unkown trend.
The El Niño phenomenon is considered unpredictable [52, 39, 6, 25, 32, 35, 50, 40]. Our analysed data are the yearly means in the The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. [^5] The data from https://psl.noaa.gov/data/timeseries/month/data/nino4.long.anom.data were retrieved on January 29th, 2026. There were no values for the last four months of 2025. The full data contains $n=155$ yearly means between 1870 and 2024.
5.1 Trend
We detect significant signals only from the weighted data. Therefore, the trend and the forecasts are solved from the weighted data. The weighted DCM analysis results for all data are given in Table S1. For one, two, three and four signals, the best trend alternative is the linear trend $K_{1}=1$. We use this $K_{1}=1$ trend in our DCM analysis of El Niño data. For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$during $\Delta T=154$ years (Equation 5), like in all $\mathcal{M}$=1-4 models.
5.2 Forecasting data
We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table S2. The periods of the three strongest signals are $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signals models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
5.3 All data
The DCM analysis results for all weighted data are given in Table S3. The three signals $\mathcal{M}$=3 model is the best. The four signal model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$. The three and four signal models are shown in Figures S3 and S4. For both alternatives, the predicted values, open circles, after 2024 are given in Table S4. Both alternatives give essentially the same forecast.
For the sake of consistency, we also give the results of non-weighted analysis of all data (Table S5). We detect the same signals as from non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower. This confirms that the data error information is crucial.
5.4 Main result
Our DCM can model and forecast the El Niño phenomenon. The solar forcing is the only possible explanation for these results. Within their error limits, the detected three El Niño signals could represent the Hale cycle [24], the solar cycle[46, 56] and half the solar cycle. These strict El Niño periodicities indicate that the planets could cause a stationary and deterministic solar cycle [28]. A stochastic and non-stationary solar cycle [8, 54] could not cause the detected regular El Niño periodicities.
Acknowledgements
This work has made use of NASA’s Astrophysics Data System (ADS) services.
Author contributions statement
L.J. peformed this research and wrote the manuscript.
Consent for publication
The author grants permission for publication.
Additional information
Supplementary information The online version contains supplementary material available.
Data availability
Main manuscript analysis: All data files and all DCM control files are stored to
@@ Acknowledgements
Appendix analysis: All files, DCM Python code, DCM control files and Python code are stored to the
https://zenodo.org/records/18595333.